OpenStudy (emilyjones284):

When comparing two jobs, Chris noticed that the hours every week were the same at each, but the job at the hospital paid $3.00 per hour more than the job at the mall. She can earn $440 per week at the hospital and $320 per week at the mall. What is the hourly wage at the mall?

OpenStudy (anonymous):

m for mall per hour h for hospital per hour m+3=h this saying chris makes 3 more dollars an hour at the h. with me so far?

OpenStudy (emilyjones284):

yes im with you

OpenStudy (anonymous):

well money per week is hourly wage times hours since both work same amount of hours we'll say c for constant. m*c= 320 at mall a week h*c=440 at hospital m+3=h from above and previous post. Now this is your system of equations you must solve to find how much he makes :) can you solve it? try :)

OpenStudy (anonymous):

I am impressed by this problem sort of. And you'll see why :D

OpenStudy (emilyjones284):

im not completley sure how to solve that haha

OpenStudy (anonymous):

do you know system of equations? substitution method or elimination method?

OpenStudy (anonymous):

Well lets try this but the top 2 equations in terms of c m*c= 320 m=320/c h*c=440 h=440/c m+3=h Now that we know what m and h equal we plug into third equation (320/c)+3=(440/c) We know solve for c Just solve for c in this equation and tell me what it is. Can you do it?

OpenStudy (anonymous):

*put not but lol

OpenStudy (emilyjones284):

would it be $8 per hour?

OpenStudy (anonymous):

YES lol who told you :P

OpenStudy (anonymous):

we then use c =40 to plug into the equation and solve for m m*40= 320

OpenStudy (anonymous):

m=320/40= 8 lol

OpenStudy (emilyjones284):

yay haha thanks (: